Hi
A packet of sweets is shared between Tom Tim and team in the ratio of 3:4:8.after Tom ate 15 sweets the ratio became 2:6:12.how many sweets did Tom have left.
Thanks

Answer by ikleyn(52754)

.

The original ratio 3:4:8 is the same as 9:12:24.


The next ratio 2:6:12 is the same as 4:12:24.


    (Notice that in both these two extended ratios the last two terms are identical (!) )


So, the beginning numbers of sweets were  9x, 12x and 24x.


After Tom ate 15 sweets, his 9x became 4x.


It means that  9x - 4x = 15,  or  5x = 15, i.e.  x = 15/5 = 3.


So, Tom left  4x = 4*3 = 12 sweets.    ANSWER

Solved.